A bullet is fired at a plank of wood with a speed of 200m/s. After passing through the plank, its speed reduces to 180m/s. Another bullet, of the same mass and size but moving with a speed of 100m/s is fired at the same plank. What would be the speed of this bullet after passing through the plank?? Assume that the resistance offered by the plank is the same for both the bullets.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

let the retardation is constant(a) then we can use

v2 = u2 +  2as

v2-u2 = 2as                         ( s is width of plank)

for first bullet , v = 180 , u = 200

1802 - 2202 = 2as      ...............1

for second bullet , v=? , u = 100 so

v2 - 1002 = 2as          ..............2              (width of plank & resistance remains same)

divide 1/2

1802 - 2002/v2-1002 = 1

v2 = 2400

v = 49m/s

Aiswarya Ram Gupta
13 years ago

thnku